OpenStudy (volleyballlover55):

What is the rational number equivalent to 3 point 24 with a bar over 24? 3 and 6 over 97 3 and 28 over 99 3 and 8 over 33 3 and 10 over 32

OpenStudy (faiqraees):

x = 3.242424242... 100x = 324.2424242424... 100x- x =324.2424242424...-3.242424242... 99x = 321 x = 321/99 x = 3 8/33

OpenStudy (volleyballlover55):

ok thanks! :)